What are the dimensions of a rectangle if the length is 5 less than three times the width and the area is 78?!

Answer by richard1234(7193)

Let W be the width, then the length is L = 3W-5

W(3W-5) = 78

Solve by factoring, quadratic formula, or noting that 78 = 6*13 and 13 = 3(6) - 5. In any case, W = 6, L = 13.